OpenStudy (anonymous):

A law firm has eight male and six female lawyers. A committee of five, two male lawyers and three female lawyers, is to be chosen to represent the firm at a conference. How many committees are possible?

OpenStudy (aric200):

560 I believe hold on

OpenStudy (anonymous):

the no of ways of chosing 2 male from 8 is C(8,2)= 8*7/2 =28 and that of chosing 3 from 6 female lawyer is C(6,3)= 6*5*4/6= 20 hence reqd no. of ways is 28*20

OpenStudy (anonymous):

:-) Thanks again Matricked for showing me the steps. I get so confused on word problems.